OpenStudy (anonymous):

solve (x-x^2)^2 plssss

OpenStudy (whpalmer4):

There's nothing there to solve. Do you mean expand it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mmm loooks hard

OpenStudy (anonymous):

lke multiply its square

OpenStudy (anonymous):

dont know if i can do it

OpenStudy (anonymous):

helpppppppp

OpenStudy (anonymous):

multiply its square

OpenStudy (anonymous):

is it (x-x^2)(X-X^2)

OpenStudy (anonymous):

is it x^4-2x^3+x^2

OpenStudy (whpalmer4):

\[(x-x^2)(x-x^2) = x(x-x^2) - x^2(x-x^2)\]by the distributive property. Now do that again and collect like terms.

OpenStudy (anonymous):

\[\left( a-b \right)^2=a^2+b^2-2ab\] \[\left( x^2 \right)^2=\left( x \right)^{2*2}=x ^{4}\]

OpenStudy (anonymous):

Are you familiar with the acronym FOIL? First, Outside, Inside, Last?

OpenStudy (whpalmer4):

FOIL is a bad crutch — useless as soon as you start multiplying anything other than binomials. Better to think of it as distributive property, repeatedly applied.

OpenStudy (whpalmer4):

\[ x^4-2x^3+x^2\] is correct, good job!

OpenStudy (anonymous):

i reather distribute

OpenStudy (anonymous):

@whpalmer4 Good point.

OpenStudy (anonymous):

wait now

OpenStudy (anonymous):

antiderive that

OpenStudy (anonymous):

pls

OpenStudy (whpalmer4):

Now you want to integrate it?

OpenStudy (anonymous):

anti

OpenStudy (whpalmer4):

the antiderivative is the integral. do you want the integral, or the derivative?

OpenStudy (anonymous):

1/5x^5-2x^4+1/3x^3 what i got

OpenStudy (anonymous):

n yes integral

OpenStudy (whpalmer4):

Well, this is an indefinite integral, so you forgot the arbitrary constant of integration. Not optional!

OpenStudy (anonymous):

pi interal from 0 to 1 the function

OpenStudy (anonymous):

antiderrive the function

OpenStudy (whpalmer4):

But we'll pretend you had it there, and check answer by taking derivative: \[\frac{d}{dx}[\frac{1}5 x^5-2x^4+\frac{1}{3}x^3+C] = 5*\frac{1}{5}x^{5-1}-4*2x^{4-1}+3*\frac{1}{3}x^{3-1}+0 \]\[= x^4-8x^3+x^2\]That's not correct...

OpenStudy (anonymous):

you know to much swag. just checking though

OpenStudy (anonymous):

umm the function is x^4-2x^3+x^2. sorry if i sound rude

OpenStudy (anonymous):

i want to antiderrive

OpenStudy (whpalmer4):

\[\int x^4-2x^3+x^2 \,dx = \frac{1}{4+1}x^{4+1}-\frac{2}{3+1}x^{3+1} + \frac{1}{2+1}x^{2+1} + C \]\[=\frac{1}{5}x^5-\frac{1}{2}x^4+\frac{1}{3}x^3+C\] Yes, and I was pointing out that you made a mistake, because when we take the derivative of what we get when we take the antiderivative, we should end up with the same thing!

OpenStudy (anonymous):

so (1/5)x^5-2x^4+(1/3)x^3 should

OpenStudy (anonymous):

Break it up into separate integrals and take the constants outside the integrands.

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

ok1/2

OpenStudy (anonymous):

well then that was rathe silly of me. all this time lol

OpenStudy (whpalmer4):

When in doubt, check your work. When not in doubt, check it anyhow :-)

OpenStudy (anonymous):

ok now multiply all of that by pi and plug 1 into x

OpenStudy (anonymous):

go i race u

OpenStudy (whpalmer4):

now hold on, you're evaluating a definite integral now? what are the limits?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

0 to 1

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

\[\frac{\pi}{30}\]

OpenStudy (anonymous):

rady

OpenStudy (anonymous):

what

OpenStudy (anonymous):

how what when where

OpenStudy (whpalmer4):

You said evaluate it from x = 0 to x = 1 and multiply by 30. I did. That's my answer.

OpenStudy (anonymous):

u go the right answer

OpenStudy (whpalmer4):

of course I did! :-)

OpenStudy (anonymous):

ok dont u need a comon denominator for the number in ()

OpenStudy (anonymous):

30

OpenStudy (anonymous):

lol just did the math. damm ur fast

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

im also stuck on another problem which i have no idea how to do start or end

OpenStudy (whpalmer4):

\[\pi \int_0^1 (x^4-2x^3+x^2)\,dx = \pi((\frac{1}{5}(1)^5 - \frac{1}{2}(1)^4+\frac{1}{3}(1)^3)- 0)= \pi(\frac{1}{5}-\frac{1}{2}+\frac{1}{3 } )\]\[= \pi\frac{6+10-15}{30} = \frac{\pi}{30}\]

OpenStudy (anonymous):

yea i got it but thanks

OpenStudy (anonymous):

can i just give you the question???

OpenStudy (anonymous):

its similar but with trig

OpenStudy (anonymous):

\[\pi \int\limits_{0}^{1}(\tan(\pi/4)x)^2dx\]

OpenStudy (anonymous):

so how do we square the tan????????

OpenStudy (whpalmer4):

Isn't that supposed to be \[\pi \int_0^1 (\tan(\frac{\pi}{4}x))^2\,dx\]

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

oh, probably take \[u = \frac{\pi}{4}x\] as the substitution, and rewrite \(\tan^2 u\) as \(\sec^2u -1\)

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

give it a try, I'll stick around and help if needed

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is sec^2u-1 the antiderrivative or just rewritten?

OpenStudy (anonymous):

???

OpenStudy (whpalmer4):

Yes?

OpenStudy (anonymous):

rewritten?

OpenStudy (whpalmer4):

I'm a little unclear on the question you are asking me here :-)

OpenStudy (anonymous):

is sec^2u-1 the antiderrivative or just rewritten?

OpenStudy (anonymous):

could you do it im kinda tired.not to be mean or anything :)

OpenStudy (anonymous):

it is the identity you use to find the anti derivative since the anti derivative of \(\sec^2(x)\) is known to be \(\tan(x)\) and the anti derivative of \(-1\) is \(-x\)

OpenStudy (anonymous):

satellite meet whpalmer, whpalmer meet satellite . you guys are smart lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if it is \(\tan^2(\frac{\pi}{4}x)\) make the substitution \[u=\frac{\pi}{4}x\] and change the limits of integration while you are at it

OpenStudy (anonymous):

\[\frac{4}{\pi}\times \pi\int_{0}^{\frac{\pi}{4}}(\sec^2(u)-1)du\]

OpenStudy (anonymous):

why did you do that???

OpenStudy (anonymous):

change limits

OpenStudy (anonymous):

why? so you don't have to go back and replace \(u\) with \(\frac{\pi}{4}x\) just stick with the \(u\)'s

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

you get \[4\tan(u)-u\] which you then evaluate at \(\frac{\pi}{4}\) and at \(0\) except you can ignore the zero part

OpenStudy (anonymous):

4-pi

OpenStudy (whpalmer4):

correctamundo...

OpenStudy (anonymous):

you found the sound

OpenStudy (anonymous):

yes indeed i did

OpenStudy (anonymous):

ok guys thanks for all your help ill be back b4 i take my test next week

OpenStudy (anonymous):

thank you <3

OpenStudy (whpalmer4):

If I don't see you before then, best of luck on the test!